Please help me solve this equation: The difference of nine times a number and 6 results in five times the difference of twice the number and 3
Answer by macston(5194)

.
x=the number
.
The difference of nine times a number and 6: 9x-6
.
Five times the difference of twice the number and 3: 5(2x-3)
.
9x-6=5(2x-3)
9x-6=10x-15 . Subtract 9x from each side.
-6=x-15 . Add 15 to each side.
9=x
.
ANSWER: x=9
.
CHECK:
9x-6=5(2x-3)
9(9)-6=5(2(9)-3)
81-6=5(18-3)
75=5(15)
75=75
.